\begin{thebibliography}{10}

\bibitem{AbernethyHaRa12}
Jacob~D Abernethy, Elad Hazan, and Alexander Rakhlin.
\newblock Interior-point methods for full-information and bandit online
  learning.
\newblock {\em IEEE Transactions on Information Theory}, 58(7):4164--4175,
  2012.

\bibitem{agarwal2017corralling}
Alekh Agarwal, Haipeng Luo, Behnam Neyshabur, and Robert~E Schapire.
\newblock Corralling a band of bandit algorithms.
\newblock In {\em Conference on Learning Theory}, pages 12--38, 2017.

\bibitem{agarwal2005efficient}
Amit Agarwal and Elad Hazan.
\newblock Efficient algorithms for online game playing and universal portfolio
  management.
\newblock {\em Electronic Colloquium on Computational Complexity, TR06-033},
  2005.

\bibitem{agarwal2006algorithms}
Amit Agarwal, Elad Hazan, Satyen Kale, and Robert~E Schapire.
\newblock Algorithms for portfolio management based on the newton method.
\newblock In {\em Proceedings of the 23rd international conference on Machine
  learning}, pages 9--16, 2006.

\bibitem{boyd2004convex}
Stephen Boyd and Lieven Vandenberghe.
\newblock {\em Convex optimization}.
\newblock Cambridge university press, 2004.

\bibitem{bubeck2018sparsity}
S{\'{e}}bastien Bubeck, Michael~B. Cohen, and Yuanzhi Li.
\newblock Sparsity, variance and curvature in multi-armed bandits.
\newblock In {\em International Conference on Algorithmic Learning Theory},
  2018.

\bibitem{bubeck2015sampling}
S{\'e}bastien Bubeck, Ronen Eldan, and Joseph Lehec.
\newblock Sampling from a log-concave distribution with projected langevin
  monte carlo.
\newblock {\em arXiv preprint arXiv:1507.02564}, 2015.

\bibitem{cover1991universal}
Thomas~M Cover.
\newblock Universal portfolios.
\newblock {\em Mathematical Finance}, 1(1):1--29, 1991.

\bibitem{cover1996universal}
Thomas~M Cover.
\newblock Universal data compression and portfolio selection.
\newblock In {\em Foundations of Computer Science, 1996. Proceedings., 37th
  Annual Symposium on}, pages 534--538. IEEE, 1996.

\bibitem{foster2016learning}
Dylan~J Foster, Zhiyuan Li, Thodoris Lykouris, Karthik Sridharan, and Eva
  Tardos.
\newblock Learning in games: Robustness of fast convergence.
\newblock In {\em Advances in Neural Information Processing Systems}, pages
  4734--4742, 2016.

\bibitem{FreundSc97}
Yoav Freund and Robert~E. Schapire.
\newblock A decision-theoretic generalization of on-line learning and an
  application to boosting.
\newblock {\em Journal of Computer and System Sciences}, 55(1):119--139, August
  1997.

\bibitem{hazan2007logarithmic}
Elad Hazan, Amit Agarwal, and Satyen Kale.
\newblock Logarithmic regret algorithms for online convex optimization.
\newblock {\em Machine Learning}, 69(2-3):169--192, 2007.

\bibitem{hazan2016introduction}
Elad Hazan et~al.
\newblock Introduction to online convex optimization.
\newblock {\em Foundations and Trends{\textregistered} in Optimization},
  2(3-4):157--325, 2016.

\bibitem{helmbold1998line}
David~P Helmbold, Robert~E Schapire, Yoram Singer, and Manfred~K Warmuth.
\newblock On-line portfolio selection using multiplicative updates.
\newblock {\em Mathematical Finance}, 8(4):325--347, 1998.

\bibitem{kalai2002efficient}
Adam Kalai and Santosh Vempala.
\newblock Efficient algorithms for universal portfolios.
\newblock {\em Journal of Machine Learning Research}, 3(Nov):423--440, 2002.

\bibitem{lovasz2006fast}
L{\'a}szl{\'o} Lov{\'a}sz and Santosh Vempala.
\newblock Fast algorithms for logconcave functions: Sampling, rounding,
  integration and optimization.
\newblock In {\em Foundations of Computer Science, 2006. FOCS'06. 47th Annual
  IEEE Symposium on}, pages 57--68. IEEE, 2006.

\bibitem{narayanan2010random}
Hariharan Narayanan and Alexander Rakhlin.
\newblock Random walk approach to regret minimization.
\newblock In {\em Advances in Neural Information Processing Systems}, pages
  1777--1785, 2010.

\bibitem{nesterov1994interior}
Yurii Nesterov and Arkadii Nemirovskii.
\newblock {\em Interior-point polynomial algorithms in convex programming},
  volume~13.
\newblock Siam, 1994.

\bibitem{orseau2017soft}
Laurent Orseau, Tor Lattimore, and Shane Legg.
\newblock Soft-bayes: Prod for mixtures of experts with log-loss.
\newblock In {\em International Conference on Algorithmic Learning Theory},
  pages 372--399, 2017.

\bibitem{Tim2018conjecture}
Tim van Erven, Dirk van~der Hoeven, and Wouter Koolen.
\newblock personal communication, 2018.

\bibitem{wei2018more}
Chen-Yu Wei and Haipeng Luo.
\newblock More adaptive algorithms for adversarial bandits.
\newblock In {\em Conference on Learning Theory}, 2018.

\bibitem{zinkevich2003online}
Martin Zinkevich.
\newblock Online convex programming and generalized infinitesimal gradient
  ascent.
\newblock In {\em Proceedings of the 20th International Conference on Machine
  Learning}, pages 928--936, 2003.

\end{thebibliography}

\documentclass{article}

\PassOptionsToPackage{numbers, compress}{natbib}

\usepackage[final]{nips_2018}

\usepackage[utf8]{inputenc} 
\usepackage[T1]{fontenc}    
\usepackage{hyperref}       
\usepackage{url}            
\usepackage{booktabs}       
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{microtype}      
\usepackage{faktor}
\usepackage{amsthm} 
\usepackage[vlined,linesnumbered,ruled]{algorithm2e}
\usepackage{bbm}
\usepackage{amsfonts}
\usepackage{amsmath}           
\usepackage{mathtools}    
\usepackage{amsopn}
\usepackage{amssymb}

\usepackage[]{color-edits}
\usepackage{soul}
\addauthor{HL}{blue}

\usepackage{scrwfile}
\TOCclone[\contentsname~(\appendixname)]{toc}{atoc}

\AfterTOCHead[toc]{%
}
\AfterTOCHead[atoc]{%
  \edef\maintocdepth{\the\value{tocdepth}}%
  \value{tocdepth}=-10000\relax%
}

\usepackage{blindtext}

\newcommand{\alg}{\textsc{BARRONS}\xspace}
\newcommand{\adaalg}{\textsc{Ada}-\textsc{BARRONS}\xspace}

\newcommand{\Reg}{\text{\rm Reg}}

\newcommand{\one}{\boldsymbol{1}}

\newcommand{\nb}{\nabla}
\newcommand{\epo}{\text{epoch}}

\DeclarePairedDelimiter\bigabs{\big\lvert}{\big\rvert}

\newcommand{\field}[1]{\mathbb{#1}}

\newcommand{\fR}{\field{R}}

\newcommand{\inner}[1]{ \left\langle {#1} \right\rangle }

\newcommand{\norm}[1]{\left\|{#1}\right\|}

\newcommand{\order}{\ensuremath{\mathcal{O}}}

\newcommand\numberthis{\addtocounter{equation}{1}\tag{\theequation}}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}

\allowdisplaybreaks

\DeclareMathOperator*{\argmin}{\arg\!\min}

\title{Efficient Online Portfolio with Logarithmic Regret}

\author{
  Haipeng Luo\\
  Department of Computer Science\\
  University of Southern California\\
  \texttt{haipengl@usc.edu} \\
  \And
  Chen-Yu Wei\\
  Department of Computer Science \\
  University of Southern California \\
  \texttt{chenyu.wei@usc.edu} \\
  \AND
  Kai Zheng \\
  Key Laboratory of Machine Perception, MOE, School of EECS, Peking University  \\
  Center for Data Science, Peking University, Beijing Institute of Big Data Research \\
  \texttt{zhengk92@pku.edu.cn} \\
}

\begin{document}

\maketitle

\begin{abstract}
We study the decades-old problem of online portfolio management and
propose the first algorithm with logarithmic regret that is not based on Cover's
Universal Portfolio algorithm and admits much faster implementation.
Specifically Universal Portfolio enjoys optimal regret $\order(N\ln T)$ for $N$ financial instruments over $T$ rounds,
but requires log-concave sampling and has a large polynomial running time.
Our algorithm, on the other hand, ensures a slightly larger but still logarithmic regret of $\order(N^2(\ln T)^4)$,
and is based on the well-studied Online Mirror Descent framework with a novel regularizer that can be implemented 
via standard optimization methods in time $\order(TN^{2.5})$ per round.
The regret of all other existing works is either polynomial in $T$ or has a potentially unbounded factor
such as the inverse of the smallest price relative.

\end{abstract}

\section{Introduction}

We consider the well-known online portfolio management problem~\citep{cover1991universal},
where a learner has to sequentially decide how to allocate her wealth 
over a set of $N$ financial instruments in order to maximize her return,
importantly under no assumptions at all on how the market behaves.
Specifically, for each trading period $t = 1, \ldots, T$, the learner first decides
the proportion of her wealth to invest on each stock,
and then by the end of the period observes the return of each stock
and continues to invest with her total wealth.
The goal of the learner is to maximize the ratio between her total wealth after 
$T$ rounds and the total wealth of the best {\it constant-rebalanced portfolio} (CRP)
which always rebalances the wealth to ensure a fixed proportion of investment for each stock.
Equivalently, the learner aims to minimize her {\it regret},
which is the negative logarithm of the aforementioned ratio.

The minimax optimal regret for this problem is $\order(N\ln T)$,
achieved by Cover's Universal Portfolio algorithm~\citep{cover1991universal}.
This algorithm requires sampling from a log-concave distribution
and all known efficient implementations have large polynomial (in $N$ and $T$) running time,
such as $\order(T^{14}N^4)$~\citep{kalai2002efficient}.\footnote{%
Recent improvements on log-concave sampling such as~\citep{bubeck2015sampling, lovasz2006fast, narayanan2010random}
may lead to improved running time, but it is still a large polynomial.
}

Online Newton Step (ONS)~\citep{hazan2007logarithmic}, on the other hand,
follows the well-studied framework of Online Mirror Descent (OMD) with a simple time-varying regularizer
and admits much faster implementation via standard optimization methods.
The regret of ONS is $\order(GN\ln T)$ where $G$ is the largest gradient $\ell_\infty$-norm encountered over $T$ rounds 
(formally defined in Section~\ref{subsec:setup})
and can be arbitrarily large making the bound meaningless.
A typical way to prevent unbounded gradient is to mix the output of ONS with a small amount of uniform distribution,
which after optimal trade-off can at best lead to a regret bound of $\order(N\sqrt{T\ln T})$.
An earlier work~\citep{agarwal2005efficient} achieves a worse regret bound of $\order(G^2N\ln(NT))$ via
an efficient algorithm based on another well-known framework Follow-the-Regularized-Leader (FTRL).

There are also extremely efficient approaches with time complexity $\order(N)$ or $\order(N\ln N)$ per round,
such as exponentiated gradient~\citep{helmbold1998line}, online gradient descent~\citep{zinkevich2003online}, 
and Soft-Bayes from recent work of~\citep{orseau2017soft}.
The first two achieve regret of order $\order(G\sqrt{T\ln N})$ and $\order(G\sqrt{T})$ respectively\footnote{%
For online gradient descent, $G$ is the largest gradient $\ell_2$-norm.
} 
while the last one achieves $\order(\sqrt{NT\ln N})$ without the dependence on $G$.
Despite being highly efficient, all of these approaches fail to achieve the optimal logarithmic dependence on $T$ for the regret.

As one can see, earlier works all exhibit a trade-off between regret and time complexity.
A long-standing open question is how fast an algorithm with optimal regret can be.
Specifically, are there algorithms with optimal regret and similar or even better time complexity compared to ONS?

In this work, we make a significant step toward answering this question
by proposing a simple algorithm with regret $\order(N^2(\ln T)^4)$ and time complexity $\order(TN^{2.5})$ per round.
To the best of our knowledge, this is the first algorithm with logarithmic regret (and no dependence on $G$)
that is not based on Cover's algorithm and admits fast implementation comparable to ONS and~\citep{agarwal2005efficient}.
As a comparison, we show in Table~\ref{tab:comparisons} the regret and time complexity of existing works and ours,
where for OMD/FTRL-type algorithms we do a naive calculation of the running time based on the Interior Point Method~\citep{nesterov1994interior} 
(to solve the key optimization problems involved), despite the possibility of even faster implementation.

Our algorithm is parameter-free and deterministic.
It follows the OMD framework with a novel regularizer that is a mixture of the one used in ONS
and the so-called log-barrier (a.k.a. Burg entropy)~\citep{foster2016learning, agarwal2017corralling, wei2018more}.\footnote{%
A recent work~\citep{bubeck2018sparsity} uses a mixture of the Shannon entropy and log-barrier as the regularizer for a different problem.
}
Critically, our algorithm also relies on an increasing learning rate schedule for the log-barrier similar to recent works on bandit problems~\citep{agarwal2017corralling, wei2018more},
as well as another sophisticated adaptive tuning method for the learning rate of the ONS regularizer, which resembles the standard doubling trick but requires new analysis since monotonicity does not hold for our problem.

 \renewcommand{\arraystretch}{1.3}
\begin{table}[t] 
  \centering
  \caption{Comparisons of regret and running time of different algorithms. Note that $G$ is potentially unbounded.
  For running time, we assume Interior Point Method is used to solve the involved optimization problems,
  and omit all logarithmic (in $N$ and $T$) factors.}
  \begin{tabular}{ | c | c | c | }
    \hline
    Algorithm & Regret & Time (per round) \\ 
    \hline 
    Universal Portfolio~\citep{cover1991universal, kalai2002efficient} & $N\ln T$ & $T^{14}N^4$ \\  
    \hline
    ONS~\citep{hazan2007logarithmic} & $GN\ln T$ & $N^{3.5}$\\
    \hline
    FTRL~\citep{agarwal2005efficient} & $G^2 N\ln (NT)$ & $TN^{2.5}$\\
    \hline
    EG~\citep{helmbold1998line}  & $G\sqrt{T\ln N}$ & $N$ \\
    \hline
    Soft-Bayes~\citep{orseau2017soft} & $ \sqrt{NT\ln N}$ & $N$ \\
    \hline
    \adaalg (\textbf{this work}) & $ N^2(\ln T)^4 $ & $TN^{2.5}$  \\
    \hline
  \end{tabular}
  \label{tab:comparisons}
\end{table}

\subsection{Notation and Setup}\label{subsec:setup}
The online portfolio problem fits into the well-studied online learning framework (see for example~\citep{hazan2016introduction}).
Formally, the problem proceeds for $T$ rounds (for some $T > N$).
On each round $t = 1, \ldots, T$, the learner first decides
a distribution $x_t \in \Delta_N$ where $\Delta_N$ is the $(N-1)$-dimensional simplex.
After that the learner observes the {\it price relative vector} $r_t \in \fR_+^N$
so that her total wealth changes by a factor of $\inner{x_t, r_t}$.
Taking the negative logarithm, this corresponds to observing a loss function $f_t(x) = -\ln\inner{x_t, r_t}$ for $x\in \Delta_N$,
chosen arbitrarily by an adversary.
The regret of the learner against a CRP parameterized by $u\in \Delta_N$ is then defined as 
\[
\Reg(u) = \sum_{t=1}^T f_t(x_t) - \sum_{t=1}^T f_t(u) = -\ln\frac{\Pi_{t=1}^T \inner{x_t, r_t}}{\Pi_{t=1}^T \inner{u, r_t}},
\]
which is exactly the negative logarithm of the ratio of total wealth per dollar invested between the learner and the CRP $u$.
Our goal is to minimize the regret against the best CRP, that is, to minimize $\max_{u \in \Delta_N} \Reg(u)$.
This setup is also useful for other non-financial applications such as data compression~\citep{cover1996universal, orseau2017soft}. 

Note that the regret is invariant to the scaling of each $r_t$ and thus without loss of generality we assume $\max_{i \in [N]} r_{t,i} = 1$ for all $t$ where
we use the notation $[n]$ to represent the set $\{1, \ldots, n\}$.
It is now clear what the aforementioned largest gradient norm $G$ formally is:
$G = \max_{t\in[T]} \norm{\nabla f_t(x_t)}_\infty = \max_{t\in [T], i \in [N]} \frac{r_{t,i}}{\inner{x_t, r_t}} \leq \min\{\frac{1}{\min_{t,i} r_{t,i}}, \frac{1}{\min_{t,i}, x_{t,i}}\}$,
which in general can be unbounded.
To control its magnitude, previous works~\citep{agarwal2005efficient, agarwal2006algorithms, hazan2007logarithmic, helmbold1998line} 
either explicitly force $x_{t,i}$ to be lower bounded, which leads to worse regret,
or make the so-called no-junk-bonds assumption (that is, $\min_{t,i} r_{t,i}$ is not too small), 
which might make sense for the portfolio problem but not other applications~\citep{orseau2017soft}.
Our main technical contribution is to show how this term can be automatically canceled 
by a negative regret term obtained from the log-barrier regularizer with increasing learning rates.

\section{Barrier-Regularized Online Newton Step}
\label{sec:fixed beta}
\begin{algorithm}[t!]
 \caption{BARrier-Regularized Online Newton Step (\alg)}
\label{algorithm:base}
 \textbf{Input}: $0<\beta\leq \frac{1}{2}, 0< \eta \leq 1$ \\
 \textbf{Define}: $\bar{\Delta}_N=\{x\in \Delta_N: x_i \geq \frac{1}{NT}, \;\forall i\}$ \\
 \textbf{Initialize}: $x_1=\frac{1}{N}\one$, $A_0=N I_N$ where $\one$ is the all-one vector and $I_N$ is the $N\times N$ identity matrix. \\

\For{$t=1,2, \ldots$}{
 Predict $x_t$ and observe loss function $f_t(x) = -\ln\inner{x, r_t}$.\\
 Make updates
\begin{align}
A_{t} &= A_{t-1} + \nabla_t \nabla_t^\top  \notag \\
\eta_{t,i} &= \eta \exp\left( \max_{s\in[t]}\log_T \frac{1}{Nx_{s,i}}  \right) \label{eq:eta} \\
x_{t+1} &= \argmin_{x\in \bar{\Delta}_N} \inner{x, \nb_t} + D_{\psi_t}(x,x_t) \label{eq:OMD}
\end{align}
where $\nabla_t = \nabla f_t(x_t)$ and $\psi_t(x) = \frac{\beta}{2}\norm{x}^2_{A_t} + \sum_{i=1}^N \frac{1}{\eta_{t,i}} \ln\frac{1}{x_i}$.
}
\end{algorithm}

Recall that for a sequence of convex regularizers $\psi_t$, the outputs of Online Mirror Descent are defined by
$x_{t+1} = \argmin_{x\in \Delta_N} \inner{x, \nabla_t} + D_{\psi_t}(x, x_t)$
where $\nabla_t$ is a shorthand for $\nabla f_t(x_t)$, $D_{\psi_t}(x, y) = \psi_t(x) - \psi_t(y) - \inner{\nabla\psi_t(y), x-y}$
is the Bregman divergence associated with $\psi_t$, and we start with $x_1$ being the uniform distribution.
The intuition is that we would like $x_{t+1}$ to have small loss with respect to a linear approximation of $f_t$,
and at the same time to be close to the previous decision $x_t$ to ensure stability of the algorithm.

Although not presented in this form originally, 
Online Newton Step~\citep{hazan2007logarithmic} is an instance of OMD with $\psi_t(x) = \frac{\beta}{2}\norm{x}_{A_t}^2 = \frac{\beta}{2} x^\top A_t x$
where $A_t = A_{t-1} + \nabla_t \nabla_t^\top$ (for some $A_0$) is the gradient covariance matrix and $\beta$ is a parameter.
The analysis of~\citep{hazan2007logarithmic} shows that the regret of ONS for the portfolio problem is $\Reg(u) = \order(\frac{N\ln T}{\beta})$ 
as long as $\beta \leq \min\big\{\frac{1}{2}, \min_{s\in[T]}\frac{1}{8|(u-x_s)^\top \nb_s |}\big\}$.
Even assuming an oracle tuning, by H\"{o}lder inequality this gives $\order(GN\ln T)$ as mentioned.

To get rid of the dependence on $G$, we observe the following.
Since $\nabla_t=-\frac{r_t}{\inner{x_t,r_t}}$, its $\ell_\infty$-norm is large only when there is a stock with high reward $r_{t,i}$ while the learner puts a small weight $x_{t,i}$ on it.
However, the reason that the weight $x_{t,i}$ is small is because the learner finds it performing poorly prior to round $t$, which means that the learner had better choices and actually should have performed better than stock $i$ previously (that is, negative regret against stock $i$). 
Now as stock $i$ becomes good at time $t$ and potentially in the future, 
as long as the learner can pick up this change quickly, the overall regret should not be too large.

Similar observations were made in previous work~\cite{agarwal2017corralling, wei2018more} for different problems in the bandit setting,
where they introduced the log-barrier regularizer with increasing learning rate to explicitly ensure a large negative regret term based on the intuition above.
This motivates us to add an extra log-barrier regularizer to ONS for our problem.
Specifically, we define our regularizer to be the following mixture:  
$\psi_t(x)\triangleq\frac{\beta}{2}\norm{x}^2_{A_t} + \sum_{i=1}^N \frac{1}{\eta_{t,i}} \ln\frac{1}{x_i}$,
where $\eta_{t,i}$ is individual and time-varying learning rate.
Different from previous work, we propose a more adaptive tuning schedule for these learning rates based on Eq.~\eqref{eq:eta}
(instead of a doubling schedule~\cite{agarwal2017corralling, wei2018more}),
but the key idea is the same: increase the learning rate for a stock when its weight is small so that the algorithm learns faster
in case the stock becomes better in the future. 
Another modification is that we force the decision set to be $\bar{\Delta}_N=\{x\in \Delta_N: x_i \geq \frac{1}{NT}, \;\forall i\}$
instead of $\Delta_N$ to ensure an explicit lower bound for $x_{t,i}$.
We call this algorithm BARrier-Regularized Online Newton Step (\alg) (see Algorithm~\ref{algorithm:base}).

Under the same condition on $\beta$ as for ONS, we prove the following key theorem for \alg 
which highlights the important negative regret term obtained from the extra log-barrier regularizer.
Note that it is enough to provide a regret bound only against smooth CRP $u \in \bar{\Delta}_N$
since one can verify that the total loss of any CRP $u\in \Delta_N$ can be approximated 
by a smooth CRP in $\bar{\Delta}_N$ up to an additive constant of 2 (Lemma~\ref{lemma:uu'} in Appendix~\ref{appendix:proofs}).

\begin{theorem}
\label{theorem:main}
For any $u \in \bar{\Delta}_N$, if $\beta \leq \alpha_T(u)$ for $\alpha_t(u)\triangleq \min\big\{\frac{1}{2}, \min_{s\in[t]}\frac{1}{8|(u-x_s)^\top \nb_s |}\big\}$, 
then \alg ensures
\begin{align}
\Reg(u) \leq \mathcal{O}\left( \frac{N\ln T}{\eta} \right) + \frac{8N\ln T}{\beta} - \frac{1}{8(\ln T)\eta} \sum_{i=1}^N \max_{t\in[T]} \frac{u_i}{x_{t,i}}. \label{eq:regret_bound}
\end{align}
\end{theorem}

The second term of Eq.~\eqref{eq:regret_bound} comes from ONS while the rest comes from the log-barrier.
To see why the negative term is useful, for a moment assume that we were able to pick $\beta$ such that 
$\frac{1}{2}\alpha_T(u^*)\leq \beta \leq \alpha_T(u^*)$ where $u^*$ is the best (smoothed) CRP.
Then by setting $\eta = \frac{1}{1024N(\ln T)^2}$, the regret against $u^*$ can be upper bounded by 
\begin{align*}
&\mathcal{O}\left( \frac{N\ln T}{\eta} \right) + \frac{16N\ln T}{\alpha_T(u^*)}- \frac{1}{8(\ln T)\eta} \sum_{i=1}^N \max_{t\in[T]} \frac{u^*_i}{x_{t,i}}\\ 
\leq & ~ \mathcal{O}\left( \frac{N\ln T}{\eta} \right) + 128N(\ln T)\max_{t\in[T]} \Bigg|\frac{\inner{r_t, u^*-x_t}}{\inner{r_t, x_t}}\Bigg|+32N\ln T - \frac{1}{8(\ln T)\eta} \sum_{i=1}^N \max_{t\in[T]} \frac{u^*_i}{x_{t,i}}\\ 
\leq & ~ \mathcal{O}\left(\frac{N\ln T}{\eta}\right)+128N(\ln T)\left(\max_{t\in[T],i}\frac{u^*_i}{x_{t,i}} + 1 \right)+32N\ln T-128N(\ln T)\sum_{i=1}^N \max_{t\in[T]} \frac{u^*_i}{x_{t,i}}\\
\leq & ~ \mathcal{O}\left(N^2(\ln T)^3\right), 
\end{align*}

which completely eliminates the dependence on the largest gradient norm $G$! 

The problem is, of course, it is not clear at all how to tune $\beta$ in this way ahead of time.
On a closer look, it is in fact not even clear whether such $\beta$ exists
since $\alpha_T(u^*)$ depends on the sequence $x_1, \ldots, x_T$ and thus also on $\beta$ itself (see Appendix~\ref{appendix:discussions} for more discussions).
Assuming its existence, a natural idea would be to run many copies of \alg with different $\beta$
and to choose them adaptively via another online learning algorithm such as Hedge~\citep{FreundSc97}.
We are unable to analyze this method due to some technical challenges (discussed in Appendix~\ref{appendix:discussions}),
let alone the fact that it leads to much higher time complexity making the algorithm impractical.
In the next section, however, we completely resolve this issue via an adaptive tunning scheme for $\beta$.

\section{\adaalg}
\label{section:adaptive beta}
Our main idea to resolve the parameter tuning issue is based on a more involved doubling trick. 
As dicussed we would like to set $\beta$ to be roughly $\alpha_T(u_T^*)$ where $u_t^* = \min_u \sum_{s\leq t} f_s(u)$.
A standard doubling trick would suggest halving $\beta$ whenever it is larger than $\alpha_t(u_t^*)$ and then restart the algorithm.
However, since $\alpha_t(u_t^*)$ is not monotone in $t$, standard analysis of doubling trick does not work.

Fortunately, due to the special structure of our problem, we are able to analyze a slight variant of the above proposal
where we halve the value of $\beta$ whenever it is larger than $\alpha_t(u_t)$, for the {\it regularized leader} $u_t$ (defined in Eq.~\eqref{eq:u_t})
instead of the actual leader $u_t^*$.
The regularization used here to compute $u_t$ is again the log-barrier, but the purpose of using log-barrier
is simply to ensure the stability of $u_t$ as discussed later. 
In fact, $u_t$ is exactly the prediction of the FTRL approach of~\citep{agarwal2005efficient}, up to a different value of the parameter $\gamma$.
Here we only use $u_t$ to assist the tunning of $\beta$.

We call the final algorithm \adaalg (see Algorithm~\ref{alg:ada}).
Note that for notational simplicity, we reset the time index back to $1$ at the beginning of each rerun,
that is, the algorithm forgets all the previous data.

\begin{algorithm}[t]
 \caption{\adaalg}
 \label{alg:ada}
\textbf{Initialize}: $\beta=\frac{1}{2}, \eta=\frac{1}{2048 N (\ln T)^2}$, $\gamma=\frac{1}{25}$ \\
    Run \alg with parameter $\beta$ and $\eta$, where after each round $t$, if the following holds: 
    \begin{align}
       \beta > \alpha_t(u_t),
       \label{check_condition}
    \end{align}
    with $\alpha_t$ defined in Theorem~\ref{theorem:main}
    and 
    \begin{equation}\label{eq:u_t}
    u_t=\argmin_{u\in\bar{\Delta}_N} \sum_{s=1}^{t} f_s(u) + \frac{1}{\gamma} \sum_{i=1}^N \ln \frac{1}{u_i}, 
    \end{equation}
    then set $\beta\leftarrow \frac{\beta}{2}$, and rerun \alg from Line~\ref{rerun_label} with time index reset to $1$. \label{rerun_label} 
\end{algorithm}

To see why this works, suppose condition \eqref{check_condition} holds at time $t$ and triggers the restart. Then we know $\alpha_t(u_t) < \beta\leq \alpha_{t-1}(u_{t-1})$. On one hand, this implies that the condition of Theorem~\ref{theorem:main} holds at time $t-1$ for $u_{t-1}$, so the regret bound \eqref{eq:regret_bound} holds for $u_{t-1}$; 
on the other hand, this also implies that the term $\frac{N\ln T}{\beta}$ in Eq.~\eqref{eq:regret_bound} is bounded by $\frac{N\ln T}{\alpha_t(u_t)}$,
which further admits an upper bound in terms of $\max_{s\in[t],i\in[N]}\frac{u_{t,i}}{x_{s,i}}$ by the same calculation shown after Theorem~\ref{theorem:main}.
Therefore, if we can show $\max_{s\in[t],i\in[N]}\frac{u_{t,i}}{x_{s,i}} \approx \max_{s\in[t-1],i\in[N]}\frac{u_{t-1,i}}{x_{s,i}}$, 
then the same cancellation will happen which leads to small regret against $u_{t-1}$ for this period. 
It is also not hard to see that $u_{t-1}$ will have similar total loss compared to the actual best CRP $u_{t-1}^*$, leading to the desired regret bound overall.

Indeed, we show in Appendix~\ref{appendix:proofs} that both $x_t$ and $u_t$ enjoy a certain kind of stability,
which then implies the following lemma.

\begin{lemma}\label{lem:stability_a_t}
If condition~\eqref{check_condition} holds at time $t$, then $\max_{s\in[t-1],i\in[N]}\frac{u_{t-1,i}}{x_{s,i}}\geq \frac{1}{2}\max_{s\in[t],i\in[N]}\frac{u_{t,i}}{x_{s,i}}$.
\end{lemma}

Call the period between two restart an \textit{epoch} and use the notation $\epo(\beta)$ to indicate the epoch that runs with parameter $\beta$. 
We then prove the following key lemma based on the discussions above.

\begin{lemma}
\label{lemma:epoch_regret}
For any $u\in \bar{\Delta}_N$, if $\text{epoch}(\beta)$ is not the last epoch, then we have
\begin{align*}
\sum_{s\in \text{epoch}(\beta)} \left( f_s(x_s) - f_s(u) \right) \leq \mathcal{O}\left(N^2(\ln T)^3\right)-\frac{8 N\ln T}{\beta}; 
\end{align*}
otherwise, 
\begin{align*}
\sum_{s\in \text{epoch}(\beta)} \left( f_s(x_s) - f_s(u) \right) \leq \mathcal{O}\left(N^2(\ln T)^3\right)+\frac{8 N\ln T}{\beta}. 
\end{align*}
\end{lemma}

With this key lemma, we finally prove the claimed regret bound of \adaalg.
\begin{theorem}
\adaalg ensures $\Reg(u) \leq \mathcal{O}\left( N^2(\ln T)^4 \right)$ for any $u\in\Delta_N$. 
\end{theorem}
\begin{proof}
Again by Lemma~\ref{lemma:uu'} it suffices to consider $u\in\bar{\Delta}_N$.
Let the number of epochs be $B$.  When $B=1$, the bound holds trivially by Lemma~\ref{lemma:epoch_regret}. 
Otherwise, the regret is upper bounded by 
\begin{align*}
B\times \mathcal{O}\left(N^2(\ln T)^3\right) + \left( \sum_{b=1}^{B-1} \frac{-8}{2^{-b}} + \frac{8}{2^{-B}}\right)N \ln T= B\times \mathcal{O}(N^2(\ln T)^3)+16N\ln T.  
\end{align*}
Since for any $t\in [T]$ and $u\in\bar{\Delta}_N$, $\frac{1}{\alpha_t(u)}\leq \order(\max_{s\in[t], i}\frac{u_{t,i}}{x_{s,i}}) \leq \order(NT)$, 
which means $\alpha_t(u) \geq \Omega(\frac{1}{NT})$, condition~\eqref{check_condition} cannot hold after $\mathcal{O}(\ln(NT))$ epochs. 
Therefore $B=\order(\ln(NT)) = \order(\ln T)$ (since $T > N$) and the regret bound follows.
\end{proof}

\textbf{Computational complexity.} 
It is clear that the computational bottleneck of our algorithm is to solve the optimization problems defined by Eq.~\eqref{eq:OMD} and Eq.~\eqref{eq:u_t}.
Suppose we use Interior Point Method to solve these two problems. 
It takes time $\mathcal{O}\left(M\sqrt{N}\log\frac{N}{\epsilon}\right)$ to obtain $1-\epsilon$ accuracy where $M$ is the time complexity
to compute the gradient and Hessian inverse of the objective~\cite{boyd2004convex}, 
which in our case is $\mathcal{O}(N^3)$ for solving $x_t$ and $\mathcal{O}(TN^2+N^3)$ for solving $u_t$.
As $T > N$, the complexity per round is therefore $\mathcal{O}(TN^{2.5})$
ignoring logarithmic factors. 
We note that this is only a pessimistic estimation and faster implementation is highly possible, especially for solving $u_t$ (given $u_{t-1}$)
in light of the efficient implementation discussed in~\citep{AbernethyHaRa12} for similar problems. 

\section{Detailed Analysis}
\label{sec:proofs}

In this section we provide the key proofs for our results.

\paragraph{Analysis of \alg}
The proof of Theorem~\ref{theorem:main} is a direct combination of the following three lemmas,
where the first one is by standard OMD analysis and analysis from~\citep{hazan2007logarithmic}
and the proof is deferred to Appendix~\ref{appendix:proofs}.

\begin{lemma}\label{lem:first_step}
Under the condition of Theorem~\ref{theorem:main},  \alg ensures for any $u \in \Delta_N$
\[\Reg(u) \leq \sum_{t=1}^T  \left(\inner{ \nb_t,x_t-x_{t+1}}  +D_{\psi_t}(u,x_t)-D_{\psi_t}(u,x_{t+1}) - \frac{\beta}{2}\inner{\nb_t,x_t-u}^2\right).
\] 
\end{lemma}


\begin{lemma}\label{lem:negative_term}
\alg with parameters $\beta\leq \frac{1}{2}$ and $\eta\leq 1$ guarantees
\begin{align*}
\sum_{t=1}^T \left(D_{\psi_t}(u,x_t)-D_{\psi_t}(u,x_{t+1}) - \frac{\beta}{2} \inner{\nb_t, x_t-u}^2\right)
\leq \mathcal{O}\left(\frac{N\ln T}{\eta}\right) - \frac{1}{8(\ln T)\eta} \sum_{i=1}^N\max_{t\in[T]}\frac{u_i}{x_{t,i}}.
\end{align*}
\end{lemma}
\begin{proof}
Define $ \phi_t(x) = \frac{\beta}{2} \norm{x}_{A_t}^2 $, $ \varphi_t(x) = \sum_{i=1}^N \frac{1}{\eta_{t,i}} \ln\frac{1}{x_i} $. Then $\psi_t(x)=\phi_t(x)+\varphi_t(x)$ and $D_{\psi_t}=D_{\phi_t}+D_{\varphi_t}$. Note that $D_{\phi_t}(x,y)=\frac{\beta}{2}\norm{x-y}^2_{A_t}$ and $D_{\varphi_t}(x,y)=\sum_{i=1}^N \frac{1}{\eta_{t,i}}h\left(\frac{x_i}{y_i}\right)$ where $h(z) = z-1-\ln z$. 
For notation simplicity, we also define $\eta_{0,i}=\eta_{1,i}$ for all $i$. Now we have
\begin{align}
&\sum_{t=1}^T \left(D_{\psi_t}(u,x_t)-D_{\psi_t}(u,x_{t+1})\right)\leq D_{\psi_0}(u,x_1)+\sum_{t=1}^T\left(D_{\psi_{t}}(u,x_{t})-D_{\psi_{t-1}}(u,x_{t})\right) \nonumber\\
\leq & ~ D_{\psi_0}(u,x_1) + 
\frac{\beta}{2}\sum_{t=1}^T\left(\norm{u-x_{t}}_{A_{t}}^2-\norm{u-x_{t}}_{A_{t-1}}^2\right) 
+ \sum_{t=1}^T\sum_{i=1}^N\left( \frac{1}{\eta_{t,i}}-\frac{1}{\eta_{t-1,i}} \right)h\left(\frac{u_{i}}{x_{t,i}}\right) \nonumber \\
= & ~ D_{\psi_0}(u,x_1) 
+ \frac{\beta}{2}\sum_{t=1}^T\inner{\nb_t, u-x_{t}}^2 
+ \sum_{t=1}^T\sum_{i=1}^N\left( \frac{1}{\eta_{t,i}}-\frac{1}{\eta_{t-1,i}} \right)h\left(\frac{u_{i}}{x_{t,i}}\right) \nonumber \\
= & ~ \mathcal{O}\left( \beta N + \frac{N\ln T}{\eta}  \right) +  \frac{\beta}{2}\sum_{t=1}^T\inner{\nb_t, u-x_{t}}^2 
+ \sum_{t=2}^T\sum_{i=1}^N\left( \frac{1}{\eta_{t,i}}-\frac{1}{\eta_{t-1,i}} \right)h\left(\frac{u_{i}}{x_{t,i}}\right). \label{eqn:before_remaining}
\end{align}
It remains to deal with $\sum_{t=2}^T\sum_{i=1}^N\left( \frac{1}{\eta_{t,i}}-\frac{1}{\eta_{t-1,i}} \right)h\left(\frac{u_{i}}{x_{t,i}}\right)$. Fix $i$, let $t=s_1, s_2,\ldots, s_M \in [2,T]$ be the rounds where $ \eta_{t,i} \neq \eta_{t-1,i}$. Define $s_0=1$ and let $\eta^{(m)}=\eta_{s_m, i}$ and $x^{(m)}=x_{s_m,i}$ for notational convenience. Note that by definition $\eta^{(m)} = \eta \exp(\log_T \frac{1}{Nx^{(m)}}) \leq  \eta \exp(\log_T T) = \eta e$. We thus have
\begin{align*}
&\sum_{t=2}^T\left( \frac{1}{\eta_{t,i}}-\frac{1}{\eta_{t-1,i}} \right)h\left(\frac{u_{i}}{x_{t,i}}\right) 
=\sum_{m=1}^M \left( \frac{1}{\eta^{(m)}} - \frac{1}{\eta^{(m-1)}} \right) h\left(\frac{u_i}{x^{(m)}}\right)\\
=&\sum_{m=1}^M \left( \frac{1-\exp\left( \log_T\frac{x^{(m-1)}}{x^{(m)}} \right)    }{\eta^{(m)}}\right)h\left(\frac{u_i}{x^{(m)}}\right)
\leq \sum_{m=1}^M \left( \frac{- \log_T\frac{x^{(m-1)}}{x^{(m)}}}{\eta^{(m)}}\right)h\left(\frac{u_i}{x^{(m)}}\right)\\
\leq & \sum_{m=1}^M \left( - \frac{\log_2\frac{x^{(m-1)}}{x^{(m)}} \Big/ \log_2 T   }{\eta e}\right)h\left(\frac{u_i}{x^{(m)}}\right)
\leq \sum_{m=1}^M \left( - \frac{\log_2\frac{x^{(m-1)}}{x^{(m)}}   }{4(\ln T)\eta }\right)h\left(\frac{u_i}{x^{(m)}}\right).
\end{align*}
We first consider the case when $x^{(M)}\leq \min\left\{\frac{1}{2N},\frac{u_i}{2} \right\}$. Because $ x^{(M)}\leq \frac{1}{2N}=\frac{x^{(0)}}{2}$ and $x^{(m)}$ is decreasing in $m$, there must exist an $m^*\in\{1,2,\ldots,M\}$ such that $ \frac{x^{(m^*-1)}}{x^{(M)}}\geq 2  $ and $ \frac{x^{(m^*)}}{x^{(M)}}\leq 2 $. The last expression can thus be further bounded by 
\begin{align*}
&\sum_{m=m^*}^M \left( - \frac{\log_2\frac{x^{(m-1)}}{x^{(m)}}   }{4(\ln T)\eta}\right)h\left(\frac{u_i}{x^{(m)}}\right)
\leq \sum_{m=m^*}^M \left( - \frac{\log_2\frac{x^{(m-1)}}{x^{(m)}}   }{4(\ln T)\eta}\right)h\left(\frac{u_i}{2x^{(M)}}\right) \\
= & -\frac{\log_2\frac{x^{(m^*-1)}}{x^{(M)}}}{4(\ln T)\eta} h\left(\frac{u_i}{2x^{(M)}}\right)
\leq -\frac{1}{4(\ln T)\eta} h\left(\frac{u_i}{2x^{(M)}}\right) \tag{$x^{(m^*-1)} \geq 2x^{(M)}$} \\
= &-\frac{1}{4(\ln T)\eta} \left( \frac{u_i}{2x^{(M)}} -1 - \ln \left(\frac{u_i}{2x^{(M)}}\right) \right)\\
= & -\frac{1}{8(\ln T)\eta}\max_{t\in[T]}\frac{u_i}{x_{t,i}} + \mathcal{O}\left(\frac{\ln(NTu_i)}{\eta\ln T}\right)
\leq -\frac{1}{8(\ln T)\eta}\max_{t\in[T]}\frac{u_i}{x_{t,i}} + \mathcal{O}\left(\frac{1+Nu_i}{\eta}\right), 
\end{align*}
where in the first inequality we use the fact for $m\geq m^*$, $ \frac{u_i}{x^{(m)}} \geq\frac{u_i}{x^{(m^*)}} \geq  \frac{u_i}{2x^{(M)}}\geq 1$,
and that $h(y)$ is positive and increasing when $y\geq 1$. 

On the other hand, if $x^{(M)}\geq \frac{1}{2N}$ or $x^{(M)}\geq \frac{u_i}{2}$, we have $ \max_{t\in[T]} \frac{u_i}{x_{t,i}}= \frac{u_i}{x^{(M)}}\leq 2Nu_i+2$
and thus
\[
\sum_{t=2}^T\left( \frac{1}{\eta_{t,i}}-\frac{1}{\eta_{t-1,i}} \right)h\left(\frac{u_{i}}{x_{t,i}}\right) \leq 0 \leq  
\frac{1}{8(\ln T)\eta} \left(-\max_{t\in[T]}\frac{u_i}{x_{t,i}}+  2Nu_i+2 \right). 
\]
Considering both cases and Eq.~\eqref{eqn:before_remaining}, we get
\begin{align*}
&\sum_{t=1}^T \left(D_{\psi_t}(u,x_t)- D_{\psi_t}(u,x_{t+1})- \frac{\beta}{2}\sum_{t=1}^T\inner{\nb_t, u-x_{t}}^2\right)\\
\leq & ~ \mathcal{O}\left(\beta N + \frac{N\ln T}{\eta}\right) + \frac{1}{8(\ln T)\eta} \sum_{i=1}^N\left(-\max_{t\in[T]}\frac{u_i}{x_{t,i}}+  2Nu_i+2 \right) + \sum_{i=1}^N \mathcal{O}\left(\frac{1+Nu_i}{\eta}\right)\\
\leq & ~ \mathcal{O}\left(\frac{N\ln T}{\eta}\right) - \frac{1}{8(\ln T)\eta} \sum_{i=1}^N\max_{t\in[T]}\frac{u_i}{x_{t,i}},
\end{align*}
finishing the proof.
\end{proof}

\begin{lemma} 
\alg guarantees $\sum_{t=1}^T \inner{\nb_t, x_t-x_{t+1}} \leq \frac{8N\ln T}{\beta}$. 
\end{lemma}
\begin{proof}
Define $F_t(x) \triangleq \inner{x, \nb_t} + D_{\psi_t}(x,x_t)$.
Using Taylor's expansion and first order optimality of $x_{t+1}$ we have
\begin{align*}
F_t(x_t) - F_t(x_{t+1}) = & ~ \nabla F_t(x_{t+1})^\top(x_t - x_{t+1}) + \frac{1}{2}(x_t - x_{t+1})^\top\nabla^2 F_t(\xi_t)(x_t - x_{t+1}) \\
\geq & ~ \frac{1}{2}(x_t - x_{t+1})^\top\nabla^2 F_t(\xi_t)(x_t - x_{t+1}) 
=  ~ \frac{1}{2} \norm{x_t - x_{t+1}}_{\nabla^2F_t(\xi_t)}^2,
\end{align*}
where $\xi_t$ is some point that lies on the line segment joining $x_t$ and $x_{t+1}$.
On the other hand, by the definition of $F_t$, nonnegativity of Bregman divergence, and H\"{o}lder inequality, we have
\begin{align*}
F_t(x_t) - F_t(x_{t+1}) =  \inner{x_t-x_{t+1}, \nb_t} - D_{\psi_t}(x_{t+1}, x_t) 
\leq  \norm{x_t - x_{t+1}}_{\nabla^2F_t(\xi_t)} \norm{\nb_t}_{\nabla^{-2}F_t(\xi_t)}.
\end{align*}
Combining the above two inequalities we get $\norm{x_t - x_{t+1}}_{\nabla^2F_t(\xi_t)} \leqslant 2\norm{\nb_t}_{\nabla^{-2}F_t(\xi_t)}$, and thus
\begin{align*}
\inner{\nb_t, x_t - x_{t+1}} &\leq \norm{\nabla_t}_{\nabla^{-2}F_t(\xi_t)}\norm{x_t - x_{t+1}}_{\nabla^{2}F_t(\xi_t)} \leq 2\norm{\nb_t}_{\nabla^{-2}F_t(\xi_t)}^2 \\
= & ~ 2 \nb_t^T (\beta A_t + \nabla^2 \varphi_t(\xi_t))^{-1} \nabla_t \leq \frac{2}{\beta} \nabla_t^\top A_t^{-1} \nabla_t,
\end{align*}
where $\varphi_t$ is the log-barrier regularizer defined in the proof of Lemma~\ref{lem:negative_term} (whose Hessian is clearly positive semi-definite).
Using Lemma 11 in \cite{hazan2007logarithmic}, we continue with
\begin{align*}
\frac{2}{\beta} \sum_{t=1}^T \nabla_t^\top A_t^{-1} \nabla_t \leq \frac{2}{\beta}\ln \frac{|A_T|}{|A_0|} \leq \frac{2N\ln (1+T^3)}{\beta} \leq \frac{8N\ln T}{\beta},
\end{align*}
where the second inequality uses the fact $ \ln|A_0|=N\ln N$ and by AM-GM inequality $ \ln|A_T|\leq N\ln \frac{\mathbf{Tr}(A_T)}{N} \leq N\ln \left(N+\frac{\sum_{t=1}^T \norm{\nb_t}^2_2 }{N}\right)\leq N\ln(N+NT^3)$ since $\norm{\nb_t}^2_2 \leq N^2T^2 (\sum_i r_{t,i}^2)/(\sum_i r_{t,i})^2 \leq N^2T^2$.
This finishes the proof.
\end{proof}

\paragraph{Analysis of \adaalg}

To prove Lemma~\ref{lem:stability_a_t}, we make use of the following stability lemmas whose proofs are deferred to Appendix~\ref{appendix:proofs}.
\begin{lemma}
\label{stabilityu}
In \adaalg, if $\gamma \leq \frac{1}{25}$, then $1-\frac{\sqrt{\gamma}}{2} \leq \frac{u_{t+1,i}}{u_{t,i}} \leq 1 + \frac{\sqrt{\gamma}}{2} $ for all $t$ and $i$.
\end{lemma}

\begin{lemma}
\label{stabilityx}
In \adaalg, if $\eta \leq \frac{1}{300}$, then $1-\frac{\sqrt{3\eta}}{2} \leq \frac{x_{t+1,i}}{x_{t,i}} \leq 1 + \frac{\sqrt{3\eta}}{2} $ for all $t$ and $i$. 
\end{lemma}

\begin{proof}[Proof of Lemma~\ref{lem:stability_a_t}]
Denote $a_{t}\triangleq \max_{s\in[t],i\in[N]}\frac{u_{t,i}}{x_{s,i}}$. 
Suppose $a_t$ attains its $\max$ at $s=s'$ and $i=i'$ (i.e., $a_t=\frac{u_{t,i'}}{x_{s',i'}}$), then when $s' \leq t-1$, we have by Lemma~\ref{stabilityu}
$
a_{t-1}\geq \frac{u_{t-1,i'}}{x_{s',i'}} = \left(\frac{u_{t-1,i'}}{u_{t,i'}}\right)a_{t}\geq \frac{a_t}{1 + \frac{\sqrt{\gamma}}{2}} \geq \frac{1}{2}a_t;
$
when $s'=t$, we have by Lemma~\ref{stabilityu} and~\ref{stabilityx}
$
a_{t-1} \geq \frac{u_{t-1,i'}}{x_{t-1,i'}} = \left(\frac{u_{t-1,i'}}{u_{t,i'}}\right)\left(\frac{x_{t,i'}}{x_{t-1,i'}}\right)a_{t}
\geq \frac{1-\frac{\sqrt{3\eta}}{2}}{1 + \frac{\sqrt{\gamma}}{2}}a_t \geq \frac{1}{2}a_t.
$
This concludes the proof.
\end{proof}

\begin{proof}[Proof of Lemma~\ref{lemma:epoch_regret}]

Define $\Gamma_t(u) = \sum_{s=1}^t f_s(x_s)-\sum_{s=1}^t f_s(u)-\frac{1}{\gamma}\sum_{i=1}^N \ln\frac{1}{u_i}$. 
Suppose condition \eqref{check_condition} holds at some time $t$ at the end of $\epo(\beta)$ and cause the algorithm to restart. Then we know that $ \beta\leq \alpha_{t-1}(u_{t-1})$ and $\beta>\alpha_t(u_t)$. The first condition guarantees that Eq.~\eqref{eq:regret_bound}  holds for $u_{t-1}$ at time $t-1$. Also, note that $u_{t-1}$ is the maximizer of $\Gamma_{t-1}$. Together they imply for any $u\in \bar{\Delta}_N$,
\begin{align}
\Gamma_{t-1}(u) \leq \Gamma_{t-1}(u_{t-1})\leq \sum_{s=1}^{t-1} f_s(x_s) - \sum_{s=1}^{t-1} f_s(u_{t-1}) \leq \mathcal{O}\left( \frac{N\ln T}{\eta} \right) + \frac{8N\ln T}{\beta} - \frac{a_{t-1}}{8(\ln T)\eta} , \label{eq:hold for t-1}
\end{align}
where we recall the notation $a_{t}\triangleq \max_{s\in[t],i\in[N]}\frac{u_{t,i}}{x_{s,i}}$. 
The second condition implies
\begin{align}
\frac{1}{\beta}&<\frac{1}{\alpha_t(u_t)}=\max\Big\{2, \max_{s\in[t]}8|\nb_s^\top(u_t-x_s)|\Big\}=\max\Big\{2, \max_{s\in[t]}8\Big|\frac{\inner{x_s, u_t-x_s}}{\inner{x_s, r_s}}\Big|\Big\} \notag\\
&\leq \max\Big\{ 2, 8\max_{s\in[t],i\in[N]}\frac{u_{t,i}}{x_{s,i}}+8 \Big\}= 8a_t + 8 \leq 16 a_{t-1} + 8, \label{eq:a_t-1_is_large}
\end{align}
where we apply Lemma~\ref{lem:stability_a_t} for the last step.
Further combining this with Eq.~\eqref{eq:hold for t-1}, and noting that $f_t(x) - f_t(u) \leq \max_i \ln\frac{u_i}{x_i} \leq \ln (NT)$ for any $x\in\bar{\Delta}_N$, we have for any $u\in\bar{\Delta}_N$, 
\begin{align*}
    \sum_{s=1}^t f_s(x_s) - \sum_{s=1}^t f_s(u)
    &\leq \ln(NT) + \sum_{s=1}^{t-1} f_s(x_s) - \sum_{s=1}^{t-1} f_s(u)\leq \ln(NT)+\Gamma_{t-1}(u)+\frac{N\ln (NT)}{\gamma}\\
    &\leq \mathcal{O}\left(\frac{N\ln T}{\eta}\right) + \frac{8N\ln T}{\beta}-\frac{1}{8(\ln T)\eta}\left(\frac{1}{16\beta}-\frac{1}{2}\right) 
    \tag{by~\eqref{eq:hold for t-1} and~\eqref{eq:a_t-1_is_large}}\\
    &\leq \mathcal{O}\left(N^2(\ln T)^3\right) - \frac{8N\ln T}{\beta}. 
\end{align*}

For the last epoch, we can apply Theorem~\ref{theorem:main} over the entire epoch and simply discard the negative term to obtain the claimed bound.
\end{proof}

\section{Conclusions and Open Problems}
We have shown that our new algorithm \adaalg achieves logarithmic regret of $\mathcal{O}(N^2(\ln T)^4)$ for online portfolio with much faster running time compared to Universal Portfolio, the only previous algorithm with truly logarithmic regret. A natural open problem is whether it is possible to further improve either the regret (from $N^2$ to $N$) or the computational efficiency without hurting the other. It is conjectured in~\citep{Tim2018conjecture} that FTRL with log-barrier~\cite{agarwal2005efficient} (i.e. our $u_t$'s) might also achieve logarithmic regret without dependence on $G$. On the pessimistic side, it is also a conjecture that it might be impossible to have the optimal regret with $\mathcal{O}(N)$ computations per round~\citep{orseau2017soft}.

\paragraph{Acknowledgements.}
The authors would like to thank Tim van Erven for introducing the problem and to thank
Tim van Erven, Dirk van der Hoeven, and Wouter Koolen for helpful discussions throughout the projects, especially on the FTRL approach.
The work was done while KZ visited the University of Southern California. KZ gratefully acknowledges financial support from China Scholarship Council.
HL and CYW are grateful for the support of NSF Grant \#1755781.

\newpage
\bibliography{bibliography} 
\bibliographystyle{plain}

\newpage
\appendix
\section{Discussions on Tuning $\beta$}
\label{appendix:discussions}
In this section, we discuss the challenges in tuning $\beta$ via other approaches.
Recall that by the calculation shown after Theorem~\ref{theorem:main}, a $\beta$ such that
$\frac{1}{2}\alpha_T(u^*)\leq \beta \leq \alpha_T(u^*)$ where $u^*\triangleq \argmin_{u\in\bar{\Delta}_N} \sum_{t=1}^T f_t(u)$
ensures a regret bound of $\mathcal{O}(N^2(\ln T)^3)$. 
We first show the existence of such $\beta$ when the environment is \textit{oblivious}, that is, $r_1, \ldots, r_T$ are all fixed ahead of time. 
(However, we emphasize that our adaptive tuning method introduced in Section~\ref{section:adaptive beta} does not rely on the existence of such $\beta$ at all
and works even against non-oblivious environments.) 

When $r_1, r_2, \ldots, r_T$ are fixed and thus $u^*$ is also fixed, one can view $\alpha_T(u^*)$ as a (complicated) function of $\beta$.
It is not hard to see that this function is continuous: note that $x_{t+1}$ is a continuous function with respect to $ \beta, A_t, \eta_t, x_t, \nb_t$ because $x_{t+1}$ is the minimizer of a strongly convex function parameterized by these quantities. Also, $A_t, \eta_t, \nb_t$ are continuous functions of $\{x_1, \ldots, x_t\}$.\footnote{%
The fact that $\eta_t$ is continuous with respect to $\{x_1, \ldots, x_t\}$ depends on our new increasing learning rate scheme
and is not true for the scheme used in previous works~\citep{agarwal2017corralling, wei2018more} based on doubling trick.
} 
So overall, $x_{t+1}$ is a continuous function of $\{\beta, x_1, \ldots, x_t\}$. By induction, we know that $x_t$ is a continuous function of $\beta$ for all $t$. Finally, since $\alpha_T(u^*)$ continuously depends on $\{x_1, \ldots, x_T\}$, it is also a continuous function of $\beta$.  

Next note that the range of $\alpha_T(u^*)$ is $\big[\frac{1}{16NT},\frac{1}{2}\big]$ because $8|\nb_t^\top(u^*-x_t)|\leq 8\norm{\nb_t}_\infty\norm{u^*-x_t}_1\leq 16NT$. Thus by intermediate value theorem, if we vary $\beta$ from $\frac{1}{32NT}$ to $\frac{1}{2}$, there must exist a $\beta$ such that $\frac{1}{2}\alpha_T(u^*) \leq \beta\leq \alpha_T(u^*)$, which completes our argument. In fact, by $\alpha_T(u^*)$'s continuity, the set of $\beta$'s satisfying the inequality will form an interval or a union of intervals. 

Given that such $\beta$ does exist but is unknown, a natural idea is to instantiate $M$ copies of \alg's with different $\beta$'s forming a grid on $[\frac{1}{32NT}, \frac{1}{2}]$, then use Hedge to learn over these copies, which only introduces an additional $\ln M$ regret since the loss is exp-concave. If any of these $\beta$'s happens to fall into one of the intervals  described above, then the algorithm has overall regret $\mathcal{O}(N^2(\ln T)^3 + \ln M)$.

However, the challenge is to figure out how dense the grid has to be, which depends on the slope (i.e. Lipschitzness) of $\alpha_T(u^*)$ with respect to $\beta$.
The larger the slope, the denser the grid needs to be. 
Trivial analysis only shows that the Lipschitzness is exponential in $T$, which is far from satisfactory.
Also note that the running time per round of this algorithm is $(MN^{3.5})$. 
Therefore even if $M$ is polynomial in $T$ which is good for the regret, it still defeats our purpose of deriving more efficient algorithms.

\section{Omitted Proofs}
\label{appendix:proofs}

We first show that competing with smooth CRP from $\bar{\Delta}_N$ is enough.

\begin{lemma}
\label{lemma:uu'}
For any $u' \in \Delta_N$, with $u =\left(1-\frac{1}{T}\right) u' + \frac{\one}{NT} \in \bar{{\Delta}}_N$
we have
\[
\sum_{t=1}^T f_t(x_t) - \sum_{t=1}^T f_t(u') \leq \sum_{t=1}^T f_t(x_t)- \sum_{t=1}^T f_t(u) + 2.
\]
\end{lemma}
\begin{proof}
By convexity of $f_t$, we have
\begin{align*}
\sum_{t=1}^T f_t(u) - \sum_{t=1}^T f_t(u') &\leq \sum_{t=1}^T \nabla f_t(u)^\top (u-u') \leq \sum_{t=1}^T \frac{(u'-u)^\top r_t}{u^\top r_t} \\
&\leq \sum_{t=1}^T \frac{\left(\frac{u}{1-\frac{1}{T}}-u\right)^\top r_t}{u^\top r_t} = \frac{1}{1-\frac{1}{T}} \leq 2.
\end{align*}
\end{proof}

Next we provide the omitted proofs for several lemmas.

\begin{proof}[Proof of Lemma~\ref{lem:first_step}]
Note that the function $h_t(x)=e^{-2\beta f_t(x)} = \inner{x, r_t}^{2\beta}$ is concave since $0 \leq 2\beta \leq 1$. 
Therefore we have
$
h_t(u)\leq h_t(x_t)+\inner{\nabla h_t(x_t), u-x_t}. 
$
Plugging in $\nabla h_t(x)=-2\beta e^{-2\beta f_t(x)}\nabla f_t(x)$ gives
\begin{align*}
e^{-2\beta f_t(u)} \leq e^{-2\beta f_t(x_t)} \left(1-2\beta \inner{\nabla_t, u-x_t}\right), 
\end{align*}
or equivalently
\begin{align*}
f_t(u)\geq f_t(x_t)-\frac{1}{2\beta}\ln\left(1-2\beta \inner{\nabla_t, u-x_t}\right)
\end{align*}
By the condition on $\beta$ we also have $\bigabs{2\beta \inner{\nabla_t, u-x_t}}\leq \frac{1}{4}$. Using the fact $-\ln (1-z)\geq z+\frac{1}{4}z^2$ for $|z|\leq \frac{1}{4}$ gives:
\begin{align*}
f_t(x_t)-f_t(u) &\leq \inner{\nb_t, x_t-u} - \frac{\beta}{2}\inner{\nb_t,x_t-u}^2 \\
&= \inner{\nb_t, x_t-x_{t+1}} + \inner{\nb_t, x_{t+1}-u}  - \frac{\beta}{2}\inner{\nb_t, x_t-u}^2 \\
&\leq \inner{\nb_t, x_t-x_{t+1}} + D_{\psi_t}(u,x_t)-D_{\psi_t}(u,x_{t+1}) - \frac{\beta}{2}\inner{\nb_t, x_t-u}^2,
\end{align*}
where the last step follows standard OMD analysis.
More specifically, since $x_{t+1}$ is the minimizer of the function $F_t(x)\triangleq \inner{\nb_t, x}+D_{\psi_t}(x,x_t)$, by the first-order optimality condition, we have $\inner{u-x_{t+1}, \nb F_t(x_{t+1})}\geq 0$ for all $u\in\bar{\Delta}_N$. Note $\nb F_t(x_{t+1})=\nb_t+\nb\psi_t(x_{t+1})-\nb\psi_t(x_{t})$. Rearranging the condition gives $ \inner{\nb_t, x_{t+1}-u} \leq \inner{\nb \psi_t(x_{t+1})-\nb\psi_t(x_{t}), u-x_{t+1}}$. Directly using the definition of Bregman divergence, one can verify $\inner{\nb \psi_t(x_{t+1})-\nb\psi_t(x_{t}), u-x_{t+1}}=D_{\psi_t}(u,x_t)-D_{\psi_t}(u,x_{t+1})-D_{\psi_t}(x_{t+1},x_t)$, which is further bounded by $ D_{\psi_t}(u,x_t)-D_{\psi_t}(u,x_{t+1})$ by the nonnegativity of Bregman divergence. This concludes the proof.
\end{proof}

\begin{proof}[Proof of Lemma~\ref{stabilityu}]
Define $\Psi_{t}(u) = \sum_{s=1}^{t} f_s(u) + \frac{1}{\gamma} \sum_{i=1}^N \ln \frac{1}{u_i}$.
We first show that if $\norm{u_t - u_{t+1}}_{\nabla^2 \Psi_{t+1}(u_t)} \leq \frac{1}{2}$ holds, then the conclusion follows. 

Indeed, note that $\nabla^2 \Psi_{t+1}(u_t) = \sum_{s=1}^{t+1}\frac{r_s r_s^\top}{\inner{u_t,r_s}^2} + \frac{1}{\gamma} \left[\frac{1}{u_{t,i}^2}\right]_{\text{diag}} \succeq \frac{1}{\gamma} \left[\frac{1}{u_{t,i}^2}\right]_{\text{diag}}$, where $\left[\frac{1}{u_{t,i}^2}\right]_{\text{diag}}$ represents the $N$ dimensional diagonal matrix whose $i$-th diagonal element is $\frac{1}{u_{t,i}^2}$. We thus have 
\begin{align*}
 & ~ \norm{u_t - u_{t+1}}_{\frac{1}{\gamma}[1/u_{t,i}^2]_{\text{diag}}} \leq \norm{u_t - u_{t+1}}_{\nabla^2 \Psi_{t+1}(u_t)} \leq 1/2,
\end{align*}
which implies 
$\frac{(u_{t,i} - u_{t+1,i})^2}{\gamma u_{t,i}^2} \leq \frac{1}{4}$, or
$
1-\frac{\sqrt{\gamma}}{2} \leq \frac{u_{t+1,i}}{u_{t,i}} \leq 1 + \frac{\sqrt{\gamma}}{2}
$ for all $i\in[N]$. 

Next, we prove the inequality $\norm{u_t - u_{t+1}}_{\nabla^2 \Psi_{t+1}(u_t)} \leq \frac{1}{2}$. Note $u_{t+1} = \argmin_{x\in \bar{\Delta}_N} \Psi_{t+1}(x)$. If we can prove $\Psi_{t+1}(u') > \Psi_{t+1}(u_t)$ for any $u'$ that satisfies $\norm{u' - u_{t}}_{\nabla^2 \Psi_{t+1}(u_t)} = \frac{1}{2}$, then we obtain the desired inequality $\norm{u_t - u_{t+1}}_{\nabla^2 \Psi_{t+1}(u_t)} \leq \frac{1}{2}$ by the convexity of $\Psi_{t+1}$. 

By Taylor's expansion, we know there exists some $\xi$ in the line segment joining $u'$ and $u_t$, such that 
\begin{align*}
\Psi_{t+1}(u') = & ~ \Psi_{t+1}(u_t) + \nabla \Psi_{t+1}(u_t)^\top (u'-u_t) + \frac{1}{2} (u'-u_t)^\top \nabla^2 \Psi_{t+1}(\xi)(u'-u_t) \\
= & ~ \Psi_{t+1}(u_t) + \nabla f_{t+1}(u_t)^\top(u'-u_t) + \nabla \Psi_{t}(u_t)^\top (u'-u_t) + \frac{1}{2} \norm{u'-u_t}^2_{\nabla^2 \Psi_{t+1}(\xi)} \\
\geq & ~ \Psi_{t+1}(u_t) + \nabla f_{t+1}(u_t)^\top(u'-u_t) + \frac{1}{2} \norm{u'-u_t}^2_{\nabla^2 \Psi_{t+1}(\xi)}\\
\geq & ~ \Psi_{t+1}(u_t) - \norm{\nabla f_{t+1}(u_t)}_{\nabla^{-2}\Psi_{t+1}(u_t)} \norm{u'-u_t}_{\nabla^{2}\Psi_{t+1}(u_t)} + \frac{1}{2} \norm{u'-u_t}^2_{\nabla^2 \Psi_{t+1}(\xi)}\\
= & ~ \Psi_{t+1}(u_t) - \frac{1}{2}\norm{\nabla f_{t+1}(u_t)}_{\nabla^{-2}\Psi_{t+1}(u_t)} + \frac{1}{2} \norm{u'-u_t}^2_{\nabla^2 \Psi_{t+1}(\xi)} \numberthis \label{medium}
\end{align*}
where the first inequality is by the optimality of $u_t$.
As $\nabla^2 \Psi_{t+1}(u_t)  \succeq \frac{1}{\gamma} \left[\frac{1}{u_{t,i}^2}\right]_{\text{diag}}$ implies $\nabla^{-2} \Psi_{t+1}(u_t) \preceq \gamma \left[u_{t,i}^2\right]_{\text{diag}}$, we continue with
\begin{align*}
\norm{\nabla f_{t+1}(u_t)}_{\nabla^{-2}\Psi_{t+1}(u_t)}^2 \leq & ~ \norm{\nabla f_{t+1}(u_t)}_{\gamma \left[u_{t,i}^2\right]_{\text{diag}}}^2 
= \frac{\gamma r_{t+1}^\top \left[u_{t,i}^2\right]_{\text{diag}} r_{t+1}}{ \inner{u_t, r_{t+1}}^2}  
= \frac{\gamma \sum_{i=1}^N u_{t,i}^2r_{t+1,i}^2}{\inner{u_t, r_{t+1}}^2} 
\leq \gamma. \numberthis \label{smallgradient}
\end{align*}
Note $\xi$ is between $u_t$ and $u'$, so $\norm{\xi - u_{t}}_{\nabla^2 \Psi_{t+1}(u_t)} \leq \frac{1}{2}$ and thus $\frac{\xi_i}{u_{t,i}} \leq 1 + \frac{\sqrt{\gamma}}{2} \leqslant \frac{11}{10}$ according to previous discussions. Therefore, we have 
\begin{equation}
\nabla^2 \Psi_{t+1}(\xi) = \sum_{s=1}^{t+1}\frac{r_s r_s^\top}{(r_s^\top \xi)^2} + \frac{1}{\gamma} \left[\frac{1}{\xi_i^2}\right]_{\text{diag}} \succeq \frac{100}{121} \left(\sum_{s=1}^{t+1}\frac{r_s r_s^\top}{(r_s^\top u_t)^2} + \frac{1}{\gamma} \left[\frac{1}{u_{t,i}^2}\right]_{\text{diag}} \right)
= \frac{100}{121} \nabla^2 \Psi_{t+1}(u_t). \label{lowerhessian}
\end{equation}
Now combining inequalities (\ref{medium}), (\ref{smallgradient}) and (\ref{lowerhessian}), we arrive at
\begin{align*}
\Psi_{t+1}(u') \geq & ~ \Psi_{t+1}(u_t) - \frac{\sqrt{\gamma}}{2} + \frac{50}{121} \norm{u'-u_t}^2_{\nabla^2 \Psi_{t+1}(u_t)} \\
= & ~ \Psi_{t+1}(u_t) - \frac{\sqrt{\gamma}}{2} + \frac{25}{242} \\
\geq & ~ \Psi_{t+1}(u_t),
\end{align*}
which finishes the proof.
\end{proof}

\begin{proof}[Proof of Lemma~\ref{stabilityx}]
The proof is similar to the proof of Lemma~\ref{stabilityu}.
Denote $F_t(x) = \inner{x, \nb_t} + D_{\psi_t}(x,x_t)$. 
We again first prove that if $\norm{x_t - x_{t+1}}_{\nabla^2 F_t(x_t)} \leq \frac{1}{2}$, then the conclusion follows.

Note $\nabla^2 F_t(x_t) = \beta A_t + \left[\frac{1}{\eta_{t,i} x_{t,i}^2}\right]_{\text{diag}} \succeq \left[\frac{1}{\eta_{t,i}x_{t,i}^2}\right]_{\text{diag}} \succeq \left[\frac{1}{3\eta x_{t,i}^2}\right]_{\text{diag}}$, because $\eta_{t,i}\leq \eta\exp\left(\log_T(\frac{NT}{N})\right)\leq 3\eta$. Thus we have 
\begin{align*}
 & ~ \norm{x_t - x_{t+1}}_{\frac{1}{3\eta}[1/x_{t,i}^2]_{\text{diag}}} \leq \norm{x_t - x_{t+1}}_{\nabla^2 F_t(x_t)} \leq 1/2, 
\end{align*}
which implies
$\frac{(x_{t,i} - x_{t+1,i})^2}{3\eta x_{t,i}^2} \leq \frac{1}{4}$ and thus
$1-\frac{\sqrt{3\eta}}{2} \leq \frac{x_{t+1,i}}{x_{t,i}} \leq 1 + \frac{\sqrt{3\eta}}{2}$ for all $i\in[N]$. 

It remains to prove the inequality $\norm{x_t - x_{t+1}}_{\nabla^2 F_t(x_t)} \leq \frac{1}{2}$. Since $x_{t+1} = \argmin_{x\in \bar{\Delta}_N} F_{t}(x)$, if we can prove $F_t(x') > F_t(x_t)$ for all $x'$ that satisfies $\norm{x' - x_{t}}_{\nabla^2 F_t(x_t)} = \frac{1}{2}$, then we obtain the desired inequality $\norm{x_t - x_{t+1}}_{\nabla^2 F_t(x_t)} \leq \frac{1}{2}$ by the convexity of $F_t$. By Taylor's expansion, there exists some $\zeta$ on the line segment joining $x'$ and $x_t$, such that 
\begin{align*}
F_t(x') = & ~ F_t(x_t) + \nabla F_t(x_t)^\top (x'-x_t) + \frac{1}{2} (x'-x_t)^\top \nabla^2 F_t(\zeta)(x'-x_t) \\
= & ~ F_t(x_t) + \nb_t^\top(x'-x_t) + \frac{1}{2} \norm{x'-x_t}^2_{\nabla^2 F_t(\zeta)} \\
\geq & ~ F_t(x_t) - \norm{\nb_t}_{\nabla^{-2}F_t(x_t)} \norm{x'-x_t}_{\nabla^{2}F_t(x_t)} + \frac{1}{2} \norm{x'-x_t}^2_{\nabla^2 F_t(\zeta)}\\
= & ~ F_t(x_t) - \frac{1}{2}\norm{\nb_t}_{\nabla^{-2}F_t(x_t)} + \frac{1}{2} \norm{x'-x_t}^2_{\nabla^2 F_t(\zeta)}. \numberthis \label{medium2}
\end{align*}

As $\nabla^2 F_t(x_t) = \beta A_t + \left[\frac{1}{\eta_{t,i} x_{t,i}^2}\right]_{\text{diag}} \succeq \frac{1}{3\eta} \left[\frac{1}{x_{t,i}^2}\right]_{\text{diag}}$, we have $\nabla^{-2} F_t(x_t) \preceq 3\eta \left[x_{t,i}^2\right]_{\text{diag}}$. Therefore
\begin{align*}
\norm{\nb_t}_{\nabla^{-2}F_t(x_t)}^2 \leq \norm{\nb_t}_{3\eta \left[x_{t,i}^2\right]_{\text{diag}}}^2 
= \frac{3\eta r_t^\top \left[x_{t,i}^2\right]_{\text{diag}} r_t}{\inner{x_t, r_t}^2} 
= \frac{3\eta \sum_{i=1}^N x_{t,i}^2r_{t,i}^2}{\inner{x_t, r_t}^2} 
\leq 3\eta. \numberthis \label{smallgradient2}
\end{align*}
Since $\zeta$ is between $x_t$ and $x'$, we have $\norm{\zeta - x_{t}}_{\nabla^2 F_t(x_t)} < \frac{1}{2}$ and thus $\frac{\zeta_i}{x_{t,i}} \leq 1 + \frac{\sqrt{3\eta}}{2} \leq \frac{21}{20}$ according to previous discussions and the fact $\eta \leqslant \frac{1}{300}$. Therefore, we have 
\begin{equation}
\nabla^2 F_t(\zeta) = \sum_{s=1}^t\frac{r_s r_s^\top}{(r_s^\top \zeta)^2} + \left[\frac{1}{\eta_{t,i}\zeta_i^2}\right]_{\text{diag}} 
\succeq  \frac{400}{441} \left(\sum_{s=1}^t\frac{r_s r_s^\top}{(r_s^\top x_t)^2} + \left[\frac{1}{\eta_{t,i}x_{t,i}^2}\right]_{\text{diag}} \right) 
= \frac{400}{441} \nabla^2 F_t(x_t). \label{lowerhessian2}
\end{equation}
Now combining inequalities (\ref{medium2}), (\ref{smallgradient2}) and (\ref{lowerhessian2}), we get
\begin{align*}
F_t(x') \geq & ~ F_t(x_t) - \frac{\sqrt{3\eta}}{2} + \frac{200}{441} \norm{x'-x_t}^2_{\nabla^2 F_t(x_t)} \\
= & ~ F_t(x_t) - \frac{\sqrt{3\eta}}{2} + \frac{50}{441} \\
\geq & ~ F_t(u_t),
\end{align*}
which finishes the proof.
\end{proof}



\end{document}